THE Cl+H_{2} → HCl+H REACTION INDUCED BY IR + UV IRRADIATION OF Cl_{2} IN SOLID PARA-HYDROGEN

\underline{SHARON C. KETTWICH} and DAVID T. ANDERSON, Department of Chemistry, University of Wyoming, Laramie, WY 82071-3838, USA; PAUL L. RASTON, Department of Chemistry, University of Alberta, Edmonton, Alberta T6G 2G2, Canada.

Matrix isolation spectroscopy is a technique which enables highly reactive species to be trapped in a host environment and studied spectroscopically. Solid parahydrogen (pH_{2}) has been employed as a matrix host because of its interesting quantum mechanical properties and also because of its general inertness towards trapped dopants. However, in some cases pH_{2} can react with impurities yielding new product molecules and providing insight into non-classical reaction pathways. In this talk I will present the results from a series of experiments where molecular chlorine (Cl_{2}) doped pH_{2} crystals were exposed to two different irradiation schemes (namely UV only or IR + UV conditions) that gave rise to very different products. Cl_{2} doped pH_{2} crystals irradiated with 355 nm UV light produced almost exclusively (99%) isolated Cl-atom photofragments, indicating the reaction Cl+H_{2}(ν = 0, J = 0) → HCl+H is not readily occurring. Cl_{2} doped pH_{2} exposed simultaneously to 355 nm UV irradiation and broadband cw IR light yielded HCl photoproducts indicating that the following reaction is playing a significant (15%) role in the in situ photochemistry: Cl+H_{2}(ν = 1, J = 0) → HCl+H. The kinetic analysis of these experiments with two very different reaction pathways for UV only or IR +UV conditions will be presented. Further, the results of current investigations involving spin-orbit excited Cl-atoms generated using 416 nm UV photons will be discussed in order to explore the intriguing possibility of non-Born Oppenheimer reaction dynamics in the simple Cl+H_{2} reaction.